OpenStudy (anonymous):

Order of operation 82 ÷ 2(15 + (7 − 2)) − 23 Please show work

OpenStudy (espex):

PEMDAS, Parenthesis, Exponents, Multiplication, Division, Addition, Subtraction.

hero (hero):

(7-2) = 5 (15+ 5) = 20 82/2 = 41 41*20 = 820 820 - 23 = 797

OpenStudy (anonymous):

Please Excuse My Dear Aunt Sally 82 ÷ 2(15 + (7 − 2)) − 23 start with parenthasis 82 ÷ 2(15 + 5) -23 then you will work on the next set of parenthesis 82÷ 30 + 10 -23 since there are no exponents or multiplication, we will move on to division 2.73 +10 - 23 now we will do addition 12.73 - 23 finally subtraction -10.27 *disclaimer* something about this answer seems off, so don't take me literally, this is simply how you would use the order of opperations

hero (hero):

You needed to provide a disclaimer because your answer is definitely wrong

OpenStudy (anonymous):

yeah i kinda knew when i was getting decimals havent done that in FOREVER